### v1

# From tests to effect sizes: Quantifying uncertainty and statistical variability in multilingual and multitask NLP evaluation benchmarks

Abstract
In this paper, we introduce a set of resampling-based methods for quantifying uncertainty and statistical precision of evaluation metrics in multilingual and/or multitask NLP benchmarks. We show how experimental variation in performance scores arises from both model- and data-related sources, and that accounting for both of them is necessary to avoid substantially underestimating the overall variability over hypothetical replications. Using multilingual question answering, machine translation, and named entity recognition as example tasks, we also demonstrate how resampling methods are useful for computing sampling distributions for various quantities used in leaderboards such as the average/median, pairwise differences between models, and rankings.
From tests to effect sizes: Quantifying uncertainty and statistical variability in multilingual and multitask NLP evaluation benchmarks
Jonne Sälevä Brandeis University jonnesaleva@brandeis.edu          Duygu Ataman Middle East Technical University dataman@metu.edu.tr          Constantine Lignos Brandeis University lignos@brandeis.edu
1 Introduction
Over the last several years, multilingual research has undergone exponential growth within NLP, both in terms of model capabilities as well as evaluation datasets. Since no paper can evaluate on all languages, it is important to determine to what extent research findings generalize from one language and from one task to another. This is particularly true in recent years with the broad applicability of large language models (LLMs), where the goal often is to draw conclusions about the models that are not task- or language-dependent.
Despite this need, evaluation setups in multilingual NLP research tend to focus on within-language evaluation and fall into two broad categories. In the first category, the same experiments are repeated across several languages and the results analyzed separately. The inferences are typically reported as one collection, e.g. “Model A significantly outperforms model B on English–Finnish and English–Turkish but not English–German” or “Model A outperforms model B on 13 out of 43 language pairs.” The second variant, depicted in Figure 1, is the “leaderboard-type” scenario in which performance is evaluated on each of the $L$ languages after which the numbers are distilled into a single performance measure using a pre-specified aggregation function such as the arithmetic mean.
Given the aggregate scores, the main interest may then be to rank the $M$ models based on the single score or two compare differences between observed scores. The most interesting questions to practitioners often revolve around whether a given model “truly” outperforms another model or whether an observed performance difference was simply random noise that could be expected to occur under replication when comparing two equally performing models.
Figure 1: Left: Notional diagram of a $M\times L$ leaderboard consisting of M models tested on L languages and replicated R times each, yielding individual observations $x_{lr}^{(m)}$. Middle: Aggregation over $L$ languages into a single scalar per model and estimating between-language variance $\nu_{m}^{2}$. Right: Aggregation over replications to estimate within-language uncertainty $\eta_{ml}$.
From a statistical perspective, estimating this “consistency with random noise” is very similar to determining whether an estimate is statistically significant. Indeed, being 1.96 ($\approx 2$) standard errors from zero is often taken as a rough equivalent to statistical significance (Gelman, 2018). To estimate such an “effect size” requires specifying the source(s) of random variation in our experiments. Typically, this variation has been treated as arising from one of two sources: on the one hand, there is data-side variability which quantifies how differently each model would perform if a slightly different test set were used for a given task. In practice this is typically done using resampled versions of the original test set, for example using the bootstrap resampling algorithm (Efron, 1979).
From a sampling perspective, data-side variability can be interpreted as sampling error associated with the process of constructing a test set for a given task by sampling data from a larger set of all possible test sets. Another source of data-related variability that appears in leaderboards and other multi-task evaluations is the choice of evaluation tasks. When ranking is performed based on aggregated quantities, the overall performance and rankings of models will vary as a function of which tasks are chosen. This can be understood as the sampling error associated with creating a evaluation benchmark out of the larger population of all possible ones that could have been constructed.
On the other hand, there is also model-side variability that will be present even if the data set is held constant. With LLMs, decoding often involves random sampling of tokens, especially with generative tasks like question answering. By drawing multiple responses for each question, the performance will fluctuate randomly around some average. Even if the inference algorithm is deterministic (e.g. beam search), any potential finetuning will likely be nondeterministic due to, for example, random shuffling of minibatches and weight initialization. This will yield slightly different parameters from run to run, which will potentially yield different responses and different values of the performance metric. From a sampling perspective, model-side uncertainty can be seen as drawing samples from the probability distribution over possible responses defined by the model.
In this paper, we provide a new perspective on how resampling-based methods can be useful in analysis of experimental results in multilingual NLP. Our specific contributions as follows.
First, through connections to statistics, we show how experimental variance can be decomposed into between- and within-language components and how the latter further decomposes into model-based and data-based components ($\sigma_{l}$ and $\tau_{l}$) which arise naturally from the structure of the experimental data and can be used in conjunction with each other when replicating experiments.
Second, using QA, machine translation, and NER as case studies, our results demonstrate that none of these sources of variation are negligible. This suggests that only analyzing one source of error may underestimate the total variation and expose researchers to the risk of drawing incorrect conclusions about the relative merits of models. We show that between-language variance accounts for most of the variance in leaderboard-style aggregate evaluations, which suggests that downstream use of an LLM may be better modeled by resampling tasks as well as replications.
Third, we show how resampling and estimates of between and within-language variation can be used to derive uncertainty-aware estimates of “overall performance,” as well as more complex quantities such as relative rankings of models and approximate distributions of pairwise performance differences between them. All of our methods run in seconds to minutes and present no major computational bottlenecks on top of the overall inference time complexity.
We provide an implementation of our approach in a toolkit which we make freely available at www.github.com/j0ma/reuben. [^1] The name reuben is an acronym for “REsampling-based Uncertainty Bounds for Evaluating NLP.
2 Related work
[…]
Figure 2: Tree diagram showing the multilevel structure of the experimental data containing $R$ replications of $L$ languages for a single model. Between-language variance $\nu_{m}^{2}$ is computed across the averages $\mu^{(m)}_{l}$ whereas within-language variance $\eta_{ml}^{2}$ corresponds to the variance among the leaf nodes $x_{lr}^{(m)}$ of each subtree.
Significance testing in ML and NLP
Within ML, statistical significance and hypothesis testing have been prevalent since at least the 1990s (e.g. Dietterich, 1998; Dror et al., 2018; Demšar, 2006), though the emphasis has often been on “finding the right test” instead of fully modeling the available data. Within NLP, significance testing is often done using permutation tests and bootstrap-based hypothesis tests (Noreen, 1989; Koehn, 2004), which have become the de-facto standard featured in state-of-the-art evaluation toolkits (e.g. Post, 2018). Significance testing based on resampling is also often used when evaluating task-specific MT metrics against human judgments in order to gauge what difference magnitudes correspond to real differences as perceived by humans(e.g. Lo et al., 2023; Kocmi et al., 2021).
In contrast, Ulmer et al. (2022) advocate for understanding model performance variation based on multiple model training runs instead of resampling, and using this model-side variation to assess statistical significance. There is no clear consensus, though, and e.g. Bethard (2022) expressly warn against using random seeds to provide estimates of score distributions based on the argument that random seeds should be optimized like other hyperparameters. Some work in NLP has also evaluated different sources of variation in evaluation of MT metrics (Xiang et al., 2022).
Evaluation practices in NLP
There has also been a healthy debate regarding best practices of test set construction and evaluation using held-out datasets in more generally. Gorman and Bedrick (2019) and others (e.g. Kodner et al., 2023; Liu and Dorr, 2024) argue that we should randomize our splits and evaluate on several versions of our test sets to reduce the chance of false positives. Søgaard et al. (2021) instead favor using approaches based on adversarial splitting, arguing that naive randomization will underestimate the variation.
Benchmarks have also received lots of attention in the general machine learning community. An excellent survey is Dehghani et al. (2021) where the authors discuss the lifecycle of overall ML benchmarks, how they become stale as well as what is random when models are evaluated. The question of what aggregation function to use in leaderboards is also explored by Tatiana and Valentin (2021) where the authors explore using both the arithmetic, geometric and harmonic means. In particular, they show that results may be wildly different depending on what aggregation measure is used. Longjohn et al. (2025) also survey and develop resampling-based and Bayesian methods for popular computer vision benchmarks for LLMs.
Related ideas in statistics
While the original bootstrap algorithm (Efron, 1979) was intended as a frequentist estimation device, a Bayesian interpretation was provided by Rubin (1981), allowing the resulting estimates (empirical distributions, intervals etc.) to be interpreted as posterior quantities. Uncertainty quantification (i.e. standard error estimation) using bootstrap and other nonparametric methods is thoroughly reviewed by Efron (1981). On the importance of understanding variation, Gelman (2005) provides an excellent argument for the importance of understanding variance components in statistical analysis more generally. For a more specific application to linguistic research, see Vasishth and Gelman (2021). The multilevel structure we use is also related to the technique of random-effects meta-analysis, see Higgins et al. (2008).
Uncertainty quantification in NLP
There exists a substantial literature (e.g. Nikitin et al., 2024; Da et al., 2025; Chen et al., 2024, 2025; Yang et al., 2024; Ye et al., 2024; Wagner et al., 2024; Blackwell et al., 2024) on a task referred to as uncertainty quantification in NLP, where the task is to estimate some properties of an LLMs distribution over continuations, $p(x|x_{\text{prompt}})$, that predicts whether the model is able to produce the correct the answer to this question. While this task quantifies per-example uncertainty, our efforts instead focus on experimental variation. i.e. estimating the noise we would expect to see under repeated experiments and hypothetical evaluation suites.
3 Statistical background
We begin by formalizing the analysis of multilingual experimental results as a statistical inference problem related to a hierarchically structured population of experimental results.
Observations
Our experimental data consists of scores generated by separately evaluating $M$ models on $L$ languages. Individual model-level observations are represented as $L$-vectors of scores:
$\mathbf{x}^{(m)}=(x^{(m)}_{1},\dots,x^{(m)}_{L})^{\top}\in\mathbb{R}^{L}$
[…]
Sources of uncertainty
Under infinite replications, our observed values $x^{(m)}_{l}$ values will center around some mean value and exhibit a degree of fluctuation around it, representing measurement error arising from both model-side randomness (e.g. nondeterministic decoding) as well as sampling error of the test set. We can represent this as $x^{(m)}_{lr}=\mu^{(m)}_{l}+\varepsilon^{\text{repl}}_{lr}$ where $\varepsilon^{\text{repl}}_{lr}$ represents the deviation of observed scores from the language-specific mean $\mu^{(m)}_{l}$. In addition to this “within-language” variation arising from replication, there is also “between-language” variability related to what tasks are included in the leaderboard. Concretely, each $\mu^{(m)}_{l}$ also differs from the global mean performance across all tasks by some amount, i.e. $\mu^{(m)}_{l}=\bar{\mu}^{(m)}+\varepsilon^{\text{lang}}_{l}$. This lets us decompose the overall observations into $x^{(m)}_{lr}=\bar{\mu}^{(m)}+\varepsilon^{\text{lang}}_{l}+\varepsilon^{\text{repl}}_{lr}$. Taking the variance of both sides, we obtain a first-principles variance decomposition into between- and within-language components $\mathbb{V}[x^{(m)}_{lr}]=\mathbb{V}[\varepsilon^{\text{lang}}_{l}]+\mathbb{V}[\varepsilon^{\text{repl}}_{lr}]=\nu_{m}^{2}+\eta^{2}_{lm}$ where $\nu_{m}^{2}$ and $\eta^{2}_{lm}$ refer to between- and within-language variation of model $m$, respectively.
Replication
For a given task, two orthogonal sources of replication noise are apparent. Model-side randomness that arises from nondeterminism in decoding/text generation. Additional model-side randomness may arise from random parts of any training/finetuning runs, related to randomized weight initialization and batching. While less common in inference-only LLM evaluation, this is still a relevant source of variability for older non-pretrained models such as LSTMs(e.g. Reimers and Gurevych, 2017) and older pretrained models such as BERT and XLM-R. Another common source of score variability is sampling error when constructing each test set $\mathcal{D}_{l}$ (e.g. Koehn, 2004; Dehghani et al., 2021). Since we tend not to have access to true sampling distribution $p(\mathcal{D}_{l})$, this variability is typically estimated by resampling with replacement from the original data and computing an empirical variance estimate. This is the approach taken by many evaluation libraries such as sacrebleu (Post, 2018) and lm-evaluation-harness (Gao et al., 2024).
From a sampling perspective, these sources of variability can be viewed as parts of the data collection process: first, given a model $m$ and a random seed $s$, we sample random responses to all examples in our test set, i.e. $x_{ls}\sim p(x_{l})$. Finetuning can be seen as sampling $\theta^{(m)}_{ls}\sim p(\theta^{(m)}_{l})$ from the distribution of all model parameters that could be obtained. On the dataset side, we sample a set of languages to evaluate on and, for each language, a test set from a larger hypothetical population of similar data in language $l$, $\mathcal{D}_{lb}\sim p(\mathcal{D}_{l})$. This decomposes the total variance further as
$\mathbb{V}[x^{(m)}_{lr}]=\underbrace{\nu_{m}^{2}}_{\text{lang}}+\underbrace{\sigma_{ml}^{2}}_{\text{seed}}+\underbrace{\tau_{ml}^{2}}_{\text{boot}}$
In our analysis, we use both seed- and bootstrap-based resampling to yield a total of $R=SB$ replications per language. While the $B$ resamples may be much cheaper to obtain than the $S$ model re-instantiations in terms of time cost, we feel that using both is helpful to properly estimate the decomposition as well as to understand whether the sources contribute equally to the total within-language variance for each language.
Why care about variance components?
Most immediately, estimates of $\eta_{l}$ (total within-lanugage variability), $\sigma_{l}$ (model-side variability), and $\tau_{l}$ (test data variability) provide the researcher with estimates of what is driving the variation in their data. High values of $\sigma_{l}^{2}$ (model-side variability) suggest that the learned distribution over responses may have high entropy or, in the case of finetuning, that the architecture may be highly sensitive to random shuffling of batches due to e.g. small amounts of training data. On the other hand, high values of $\tau_{l}^{2}$ (test data variability) and $\nu_{m}^{2}$ (between-language variance) indicate that a model’s performance is particularly sensitive to the exact composition of a test set or benchmark and may suggest lower ability to generalize due to overfitting. Variance components are also intimately tied to statistical significance: we can, roughly speaking, judge an estimate as being statistically significant if it lies more than two standard errors from zero (e.g. Gelman, 2018).
Estimating model- and data-side SD
Using the predictions of $S$ replications on the original test data (i.e. no bootstrapping), we estimate $\sigma_{l}$ using the sample standard deviation formula.Since we have observations of each of the seeds on all $B$ datasets, we also compute an estimate of the standard error $\hat{\text{se}}(\hat{\sigma}_{l})$ using the sample standard deviation formula computed over the bootstrap datasets.
To estimate the boot-to-boot variability $\tau_{l}$, we first compute the F1 score variance over the $B$ bootstrap datasets separately for each of the $S$ seeds.We then construct the average estimator $\bar{\tau}_{l}=\sum_{s=1}^{S}\hat{\tau}_{ls}/S$. The standard error of $\hat{\tau}_{l}$ is estimated using the usual sample standard deviation formula over seeds.This gives us the plug-in estimate of the standard error of the averaged estimator $\hat{\text{se}}(\bar{\tau}_{l})=\hat{\text{se}}(\hat{\tau}_{ls})/\sqrt{S}$.
Aggregate scores and their standard errors
In most leaderboard-style scenarios, aggregated scalar performance measures such as the arithmetic mean $\bar{x}^{(m)}_{1:L}=\sum_{l}x^{(m)}_{l}/L$ tend to be used instead of the full score vectors. Its popularity of the arithmetic mean can be explained by how simple it is to compute, as well as the closed-form expression for its standard deviation $\text{sd}(x^{(m)}_{1},\dots,x^{(m)_{L}})/\sqrt{L}$ using only the between-language SD $\nu_{m}$ and the within-language SDs $\eta_{ml}$. Other aggregation functions, such as the geometric mean, or median may also be used, although they may not be as well-behaved in terms of standard error. For such aggregation functions, closed form expressions of the SD do not typically exist and must be estimated using resampling.
4 Task 1: LLM-based question answering
As our first case study, we focus on multilingual question answering and evaluate four LLMs on all subsets of using XQUAD (Artetxe et al., 2020). Specifically, we use aya-expanse-8B (Dang et al., 2024), TowerInstruct-Mistral-7B-v0.2 (Alves et al., 2024), Google’s gemma2-9b and finally Clarus-7B-v0.3. We evaluate using token-level F1 score, using the standard lm-evaluation-harness implementation. Details of the corpus, experimental settings and hardware used are in the appendix.
4.1 Variance components
The lm-eval-harness library we use to run our experiments automatically computes bootstrapped standard errors for F1. In addition to the bootstrap SE, we incorporate model-side uncertainty by sampling 5 answers for each question from each model. The summarized variance components are displayed in Table 1. Full decomposition is available in Table 10 in the appendix.
Model Mean SD Min Max
Between-language ($\nu$)
Clarus-7B 8.92 - - -
TowerInstruct-7B 11.03 - - -
aya-expanse-8B 12.01 - - -
gemma2-9B 5.05 - - -
Total within-language ($\eta_{ml}$)
Clarus-7B 0.89 0.41 0.40 1.67
TowerInstruct-7B 0.73 0.32 0.19 1.21
aya-expanse-8B 1.47 0.18 1.18 1.73
gemma2-9B 1.21 0.16 0.93 1.51
Boot-to-boot ($\tau_{ml}$)
Clarus-7B 0.70 0.29 0.24 1.27
TowerInstruct-7B 0.59 0.28 0.15 1.05
aya-expanse-8B 1.24 0.11 1.06 1.42
gemma2-9B 0.95 0.10 0.80 1.18
Seed-to-seed ($\sigma_{ml}$)
Clarus-7B 0.53 0.33 0.16 1.25
TowerInstruct-7B 0.41 0.18 0.12 0.74
aya-expanse-8B 0.76 0.30 0.25 1.23
gemma2-9B 0.74 0.20 0.48 1.09
Table 1: Summary of variance components for QA experiments.
Clarus-7B TowerInstruct-7B aya-expanse-8b
TowerInstruct-7B aya-expanse-8b gemma-2-9b aya-expanse-8b gemma-2-9b gemma-2-9b
Arabic 8.75 $\pm$ 1.13 -24.27 $\pm$ 1.34 -0.11 $\pm$ 1.50* -33.02 $\pm$ 1.39 -8.86 $\pm$ 1.53 24.16 $\pm$ 1.71
Chinese -3.96 $\pm$ 1.50 -22.80 $\pm$ 1.66 -6.34 $\pm$ 1.47 -18.83 $\pm$ 1.84 -2.38 $\pm$ 1.67* 16.45 $\pm$ 1.82
English 5.21 $\pm$ 1.67 -33.30 $\pm$ 1.86 8.79 $\pm$ 1.95 -38.50 $\pm$ 1.67 3.58 $\pm$ 1.75 42.09 $\pm$ 1.95
German -6.43 $\pm$ 1.41 -32.16 $\pm$ 1.44 -9.04 $\pm$ 1.46 -25.73 $\pm$ 1.72 -2.61 $\pm$ 1.74* 23.12 $\pm$ 1.77
Greek 16.45 $\pm$ 1.12 -20.77 $\pm$ 1.69 1.02 $\pm$ 1.58* -37.22 $\pm$ 1.59 -15.43 $\pm$ 1.47 21.79 $\pm$ 1.94
Hindi 20.43 $\pm$ 0.94 -26.75 $\pm$ 1.51 1.26 $\pm$ 1.46* -47.18 $\pm$ 1.34 -19.17 $\pm$ 1.27 28.01 $\pm$ 1.76
Romanian 2.36 $\pm$ 1.45* -30.14 $\pm$ 1.55 -3.32 $\pm$ 1.81* -32.50 $\pm$ 1.69 -5.69 $\pm$ 1.95 26.81 $\pm$ 2.01
Russian -6.20 $\pm$ 1.59 -20.23 $\pm$ 1.65 5.36 $\pm$ 1.91 -14.02 $\pm$ 1.65 11.57 $\pm$ 1.92 25.59 $\pm$ 1.96
Spanish 8.01 $\pm$ 2.05 -17.36 $\pm$ 1.98 12.55 $\pm$ 2.23 -25.37 $\pm$ 1.70 4.54 $\pm$ 1.99 29.91 $\pm$ 1.92
Thai 21.01 $\pm$ 1.23 -3.83 $\pm$ 1.83 6.42 $\pm$ 1.69 -24.84 $\pm$ 1.52 -14.60 $\pm$ 1.35 10.25 $\pm$ 1.90
Turkish 7.06 $\pm$ 0.90 -25.13 $\pm$ 1.50 -7.01 $\pm$ 1.38 -32.19 $\pm$ 1.53 -14.07 $\pm$ 1.40 18.12 $\pm$ 1.85
Vietnamese 3.23 $\pm$ 1.06 -29.35 $\pm$ 1.41 -5.82 $\pm$ 1.44 -32.59 $\pm$ 1.39 -9.05 $\pm$ 1.41 23.53 $\pm$ 1.68
Table 2: Pairwise differences between models on XQUAD. Non-significant differences are indicated with an asterisk.
4.2 Resampling for model comparison
We seek to quantify the uncertainty in the XQUAD F1 scores in each language as well as aggregated across languages. Specifically, we compute pairwise model differences on each language as well as between aggregate scores computed using the arithmetic mean, geometric mean and median. We also compute rankings as point estimates. Even for the arithmetic mean, it is not possible to do it in closed form since we have no estimate of the covariance between the models. Instead, we estimate the difference of the models using the following parametric bootstrap resampling procedure. Given a language, we first take an estimate of its each model’s average performance on it by averaging the $S=5$ observed values. We then use estimate of the within-language SD, $\eta_{ml}$, to get a randomly resampled performance score for each model. We estimate $\eta_{ml}$ by first computing the standard deviation of scores across seeds and then use the average bootstrap SD $\tau_{ml}$ across the $S=5$ values computed by lm-eval-harness to form an aggregate SD estimate of $\eta_{ml}$. We then sample a noise term using a random draw from $\varepsilon\sim\mathcal{N}(0,\eta_{ml}^{2})$ yielding a resampled score $x+\varepsilon$ with variance $\eta_{ml}^{2}=\sigma_{ml}^{2}+\tau_{ml}^{2}$. While the score itself may not be plausibly Gaussian, we feel that modeling the noise terms as such is reasonable, as we observe mostly small score deviations across both seeds and bootstrap resamples. We repeat this process $R=1000$ times. For each replication, we compute a performance difference on each language, and an aggregated score of performance scores across languages, using the arithmetic mean, geometric mean and median. We also rank the models based on the randomly replicated performance scores. The results of this procedure are summarized in several tables below. Table 2 summarizes pairwise differences between models and their associated SDs. Table 3 gives the rank distribution of each model using each of the summary statistics. While TowerInstruct and aya reliably place first and last regardless of metric, the rank distributions of Clarus-7b and gemma2-9b can vary significantly depending on which aggregation function is used.
Arithmetic mean
TI-7B Clarus-7B gemma2-9B aya-8b
1 100.00 - - -
2 - 24.50 75.50 -
3 - 75.50 24.50 -
4 - - - 100.00
Median
1 100.00 - - -
2 - 35.30 64.70 -
3 - 64.70 35.30 -
4 - - - 100.00
Geometric mean
1 100.00 - - -
2 - 94.90 5.10 -
3 - 5.10 94.90 -
4 - - - 100.00
Table 3: Distribution of model ranks on XQUAD using different aggregators. Simulation computed over 1,000 randomly sampled performance scores.
5 Task 2: LLM-based machine translation
[…]
5.1 Variance components
The lm-eval-harness library we use to run our experiments automatically computes bootstrapped standard errors for BLEU, ChrF++ and COMET. In addition to the bootstrap SE, we incorporate model-side uncertainty by sampling 5 translation hypotheses for each source sentence from each model. The summarized results are shown in Table 8 in the appendix. A detailed variance decomposition of the within-language variation is shown in Table 11 in the appendix. As with XQUAD, the between-language variability ($\nu$) is much larger than either the model-side or bootstrap uncertainty. Most variance components are also less than 1.0 BLEU or ChrF++ points.
5.2 Resampling for model comparison
Like with XQUAD, we also wish to quantify the uncertainty in the performance differences between the two models on each language. Given observed BLEU/ChrF++/COMET scores for each language we can easily compute a point estimate by subtraction. We quantify the uncertainty in this point estimate by resampling as with XQUAD. The estimated differences are displayed in Table 7. The most obvious observation is that Clarus-7B clearly underperforms relative to aya-expanse-8b and TowerInstruct-7B. The latter two models perform more similarly to each other and significant differences were found in 6 out of 16 translation using BLEU. Out of the 6 significant differences, TowerInstruct-Mistral-7B beat aya-expanse-8b 4 out of 6 times. When using ChrF++, significant differences were found in 10 out of 16 translation tasks, with TowerInstruct-Mistral-7B as the winner in all of them.
6 Task 3: NER
As a third task, we show how resampling can be used for model comparison in multilingual NER when finetuning pretrained language models. Our analysis is based on the OpenNER 1.0 multilingual NER benchmark (Palen-Michel et al., 2024) which consists of 61 unique datasets covering a total of 51 languages . Since we reanalyze the results of the original paper, we use the same models: mBERT and XLM-R. For each model architecture, we vary the training data according to two conditions on each language-dataset pair: (i) individual, where each model is finetuned with data from only a given training set and (ii) concatenated, where each model is finetuned with a concatenated and downsampled version of all the training data included in OpenNER.
In addition to the original results, we construct $B=100$ bootstrap replicated data sets and reuse the same predictions from the original test set. This gives us a total of $R=SB=1000$ replicated scores for each of the 61 datasets.
6.1 Variance components via resampling
Like with QA and MT, we investigate how resampling-based inference can be used to better make sense of experimental variation by estimating the variance components due to model- and data-side uncertainty for each of the 61 languages pairs that comprise the overall benchmark. A concise display of inferences for $\nu$, $\sigma$ and $\tau$ are shown in Table 4. A more detailed table of estimates is shown in the Appendix in Table 12.
Finetuning data
Model Individual Concatenated
Glot500 8.27 6.30
$\nu$ mBERT 9.95 14.38
XLM-R 7.13 6.31
Glot500 $1.59_{(2.51)}$ $0.61_{(0.39)}$
$\sigma_{l}$ mBERT $1.15_{(2.55)}$ $0.81_{(0.45)}$
XLM-R $1.24_{(3.53)}$ $0.64_{(0.42)}$
Glot500 $27_{(0.95)} 1.1$$17_{(0.79)}$
$\tau_{l}$ mBERT $1.36_{(0.86)}$ $1.47_{(0.91)}$
XLM-R $23_{(0.84)} 1.1$$18_{(0.79)}$
Table 4: Summary of variance components for NER. The three groups of rows, $\nu$, $\sigma$ and $\tau$ correspond to between-language, seed and bootstrap SDs. Subscripts indicate standard deviations across languages.
A few general observations are apparent. First, both seed and bootstrap variability tend to be less than 1.0 F1 points in most cases. Second, the average total within-language variability ($\eta_{l}$) is approximately 1.5 to 2 F1 points, suggesting that any deviations less than 3-4 F1 points will be statistically insignificant on any individual dataset.
When we compare models in the “individual” versus the “concatenated” conditions, there is a distinct decrease in the variability due to random seeds in the “concatenated” condition. This suggests that smaller languages may indeed benefit from larger finetuning datasets. At the same time, the variability due to bootstrapped test sets remains relatively constant, which suggests that larger finetuning sets do not help much in terms of dataset-to-dataset generalization.
6.2 Resampling for model comparison
To approximate quantities other than variance components, we resample the $R=SB=1000$ replications within each language. Unlike our QA and MT experiments, we do this entirely nonparametrically. We also show how the between-language score variability ($\nu$) can be incorporated into the analysis by also resampling what languages are considered when estimating overall performance. Our analysis based on effect sizes in Table 5 shows that underestimating variation may lead to an inflated sense of statistical significant differences between models results than when using randomization.
Average/overall performance
[…]
To estimate the standard error of each estimator, we sample $R=5000$ draws from the approximate joint distribution where we draw independently from the empirical distribution of scores for each language $l$. We conduct two separate simulations: one where the $L=61$ languages are kept constant and another one where they are randomly resampled with replacement. To compute an interval we use the Gaussian-inspired [^2] This is based on the property that approximately 95% of the variation lies within within $M=1.96\approx 2$ standard deviations of the point estimate. interval of $\text{estimate}\pm 2\text{se}$ Alternatively, we also use a percentile interval 2.5% and 97.5% percentiles of each estimator, computed over resampled draws. A third possible interval [^3] Used, for example, by the sacreBLEU package in machine translation evaluation. is to take half of the distance between the percentiles and report the interval estimate $\text{estimate}\pm\text{half-width}$.
The detailed results can be seen in Table 13 in the Appendix. The column F1 (true) indicates the observed F1 score, computed as an average over the $S$ seeds on the original data $\mathcal{D}$. All three types of CIs seem to consistently report similar ranges for the same estimator regardless of whether the languages are randomized. The bias and variance of the different estimators differ depending on the experimental condition. When the languages are kept constant, standard errors range from approximately 0.1 to 0.5 F1 points, except for XLM-R (ind.), where $\hat{\text{se}}(\hat{\mu}_{gm})=1.19$. When the languages are randomized, the standard errors increase considerably. All standard errors are now at least .5 F1 with many significantly larger. This is most likely related to increased variation in the normalization constant of the weights as the languages are randomized. Finetuning on all languages helps in most cases even with the languages randomized, both in terms of observed F1 scores and the observed variation. The exception to this is mBERT for which finetuning on all languages concatenated hurts performance.
Pairwise differences
To obtain estimates of the pairwise differences in the overall population-level means, $\Delta_{ml}=\hat{\mu}^{(m)}_{am}-\hat{\mu}^{(n)}_{am}$, we compute a pairwise difference matrix for each replication $r$. Given the $L\times L\times R$-dimensional array of pairwise differences, we average over the $R$-dimension to obtain a single $L\times L$ table of average differences. This is displayed as the top table in Table 5. We also compute a standard deviation of each pairwise difference and divide the mean by the SD to obtain an estimate of the effect size. Looking at the second part of Table 5, all effect sizes except the differences between XLM-R (ind.) and Glot500 (ind.) as well as XLM-R (conc.) and XLM-R (ind.) appear significantly different from zero.
We also conduct another simulation where we subsample $10$ datasets without replacement from the original 61 and estimate effect sizes as above. The motivation for this simulation is to incorporate between-language variability into the estimation of overall performance. Intuitively, this also corresponds to a simulating the effect of downstream usage of the tested models on smaller sets of tasks. These results are shown in the bottommost part of Table 5. Overall, the effect size estimates are much smaller, with many shrinking below magnitude 2 due to the increased standard deviation. This suggests that using fixed benchmarks may understate the differences between two models.
Mean: $\bar{\Delta}$
Model Glot500 (i) mBERT (c) mBERT (i) XLM-R (c) XLM-R (i)
Glot500 (c) 2.01 12.24 8.63 0.79 2.00
Glot500 (i) 10.23 6.63 -1.22 -0.01
mBERT (c) -3.60 -11.45 -10.24
mBERT (i) -7.84 -6.63
XLM-R (c) 1.21
Effect size: $\bar{\Delta}/\text{se}(\bar{\Delta})$
Original Glot500 (i) mBERT (c) mBERT (i) XLM-R (c) XLM-R (i)
Glot500 (c) 3.97 41.70 18.73 2.43 3.07
Glot500 (i) 23.57 10.54 -2.38 -0.17
mBERT (c) -9.99 -38.85 -19.21
mBERT (i) -16.49 -9.61
XLM-R (c) 1.78
Subsampled Glot500 (i) mBERT (c) mBERT (i) XLM-R (c) XLM-R (i)
Glot500 (c) 1.35 3.61 4.14 1.15 1.22
Glot500 (i) 2.52 3.25 -0.82 -0.00
mBERT (c) -0.73 -3.22 -2.42
mBERT (i) -4.05 -3.31
XLM-R (c) 0.81
Table 5: Means and effect sizes of pairwise F1 score differences $\Delta_{ml}=\hat{\mu}^{(m)}_{am}-\hat{\mu}^{(n)}_{am}$ based on $R=5000$ simulations.
Ranks
Glot500 mBERT XLM-R
Conc. Ind. Conc. Ind. Conc. Ind.
1 99% - - - 1% -
2 1% 1% - - 97% 1%
3 - 44% - - 2% 54%
4 - 55% - - - 45%
5 - - 1% 99% - -
6 - - 99% 1% - -
Table 6: Frequency distribution of ranks for each model based on $R=5,000$ bootstrap simulations of aggregate performance using the arithmetic mean $\hat{\mu}_{am}(\mathbf{x}^{(m)})$.
To estimate a distribution for model rankings, we rank each replication of the models based on the arithmetic mean across tasks and compute an empirical distribution approximation. The results of this simulation using are shown in Table 6. The bolded elements indicate the ranks observed using the original data. Overall, it seems like the ranks are stable for most models, except for Glot500 (ind.) and XLM-R (ind.) which compete for the 3rd place. We also repeated the simulation with fixed languages but observed that the ranks remained constant. This is in line with the observation from Section 6.1 that the between-language variation comprises by far the largest variance component. We suspect that under the “subsampling” condition, the ranks would display a lot more variability but leave this for future work.
7 Conclusion
In this paper, we have shown how replication and resampling can be helpful tools in the evaluation of multilingual NLP models. Through experiments on question answering, machine translation with LLMs, as well as NER using BERT/XLM-R-style PLMs, we showed how resampling-based inference can be useful in estimating variance components which can help researchers understand what their models are robust to and how much they might be underestimating total variation when computing standard errors. Since many hypothesis tests are based on computing standard errors, we urge researchers to estimate variance components to hedge against false positive discoveries.
We also showed how to approximate the distribution of score vectors using the parametric and nonparametric bootstrap, which allows us to estimate difficult-to-estimate parameters numerically, such as standard errors of pairwise differences or aggregate functions whose variance is unknown and rank distributions. We showed that in leaderboard-style aggregate evaluations, the between-task term may dominate the variance of the aggregate estimate.
Given these findings, we stress the importance of not relying on static leaderboards that, by construction, underestimate the total variability by neglecting the variance across tasks. We hope that this paper can stimulate further research into resampling-based evaluation in NLP that goes beyond hypothesis testing and computing error bars for individual performance scores.
Limitations
While the total number of languages we investigate is quite large, the languages are not evenly distributed across the tasks we study. Due to the length limit of this venue and the limitations of existing work, we are only able to present analysis of three tasks in this paper. We have chosen tasks with 100% human-generated annotation because we believe the results could be skewed if we include datasets generated via system output (e.g. Pan et al., 2017), including LLM-generated annotation or automatically translated datasets.
Ethical considerations
As a position paper, this paper argues for more careful evaluation of NLP experiments, particularly when multiple languages and models are studied in parallel. We believe this can be beneficial to the larger NLP community as it may help practitioners avoid drawing poorly supported conclusions and thus avoid making non-generalizable claims about research findings. That said, our methods are slightly more involved than traditional methods (e.g. hypothesis testing) and may theoretically lead to more conservative statistical analysis which may in theory obscure real effects.
Acknowledgments
This work was supported by the grant Improving Relevance and Recovery by Extracting Latent Query Structure by eBay to Brandeis University. This work was also supported by Brandeis University through internal research funds.
Appendix A Experimental settings
A.1 Question answering
We evaluate four LLMs on all XQUAD subsets: aya-expanse-8B (Dang et al., 2024), TowerInstruct-Mistral-7B-v0.2 (Alves et al., 2024), Google’s gemma2-9b and finally Clarus-7B-v0.3. All models are taken from the Open LLM Leaderboard [^4] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard from HuggingFace. Details of the experimental settings are in the Appendix. All experiments are run using the lm-evaluation-harness library (Gao et al., 2024) on A40 40GB GPUs using vLLM as an inference back-end (Kwon et al., 2023). We evaluate using token-level F1 score, using the standard lm-evaluation-harness implementation. We do not consider exact match as it is less robust to random deviations.
A.2 Machine translation
[…]
Appendix B Datasets used
B.1 XQUAD
XQUAD (Artetxe et al., 2020) is a a multilingual extension of the SQuAD v1.1 question answering benchmark (Rajpurkar et al., 2016). XQUAD includes translations 1,190 questions, originally in English, into Arabic, German, Greek, Spanish, Hindi, Romanian, Russian, Thai, Turkish, Vietnamese, and Mandarin Chinese.
B.2 OpenNER 1.0
Our analysis is based on the OpenNER 1.0 multilingual NER benchmark (Palen-Michel et al., 2024) which consists of 61 unique datasets covering a total of 51 languages. The authors experiment with three pretrained language models (PLM): mBERT (Devlin et al., 2019), XLM-R (Conneau et al., 2020), and Glot500-m (Imani et al., 2023). Each PLM is finetuned using two experimental conditions per language. In the “individual” condition, each PLM is finetuned separately on the train split of each of the 61 datasets. In the “concatenated” setting, the train splits of all 61 datasets are concatenated [^5] Due to computational considerations, before concatenation, the highest-resource datasets are downsampled. together before finetuning each of the three PLMs on it. The authors run each experiment 10 times using different seeds (values 42–51).
We use model outputs from the original authors with their permission. In addition to the original results, we construct $B=100$ bootstrap replicated data sets and reuse the same predictions from the original test set. This gives us a total of $R=SB=1000$ replicated scores for each of the 61 datasets.
Appendix C Additional Tables and Figures
Additional tables follow.
C.1 Variance components
aya-expanse-8b TowerInstruct-Mistral-7B
Task TowerInstruct-Mistral-7B Clarus-7B Clarus-7B
BLEU
de-en 0.40 $\pm$ 1.34 12.31 $\pm$ 1.14 11.91 $\pm$ 1.46
en-de -1.13 $\pm$ 1.13 12.48 $\pm$ 0.88 13.61 $\pm$ 1.03
en-es -0.35 $\pm$ 0.79 8.64 $\pm$ 0.76 8.99 $\pm$ 0.90
en-fr 1.01 $\pm$ 1.18 15.32 $\pm$ 1.32 14.30 $\pm$ 1.37
en-it 1.17 $\pm$ 0.74 11.96 $\pm$ 0.81 10.79 $\pm$ 0.80
en-ko 1.88 $\pm$ 0.49 5.15 $\pm$ 0.45 3.27 $\pm$ 0.35
en-nl 0.40 $\pm$ 0.90 10.05 $\pm$ 0.72 9.65 $\pm$ 0.80
en-pt -7.16 $\pm$ 1.03 12.85 $\pm$ 1.32 20.01 $\pm$ 1.35
en-ru -0.25 $\pm$ 0.94 12.31 $\pm$ 0.76 12.56 $\pm$ 0.82
es-en -1.29 $\pm$ 1.06 9.41 $\pm$ 0.93 10.70 $\pm$ 0.93
fr-en -1.54 $\pm$ 1.33 14.05 $\pm$ 1.13 15.59 $\pm$ 1.37
it-en 2.47 $\pm$ 1.10 11.16 $\pm$ 0.98 8.69 $\pm$ 1.21
ko-en -4.15 $\pm$ 0.96 6.51 $\pm$ 0.81 10.66 $\pm$ 0.95
nl-en -1.78 $\pm$ 1.01 10.14 $\pm$ 0.90 11.92 $\pm$ 1.09
pt-en -5.67 $\pm$ 1.07 18.38 $\pm$ 1.26 24.05 $\pm$ 1.24
ru-en -2.86 $\pm$ 0.95 10.01 $\pm$ 1.06 12.86 $\pm$ 1.04
ChrF++
de-en -2.08 $\pm$ 0.76 6.53 $\pm$ 0.94 8.62 $\pm$ 0.94
en-de -2.03 $\pm$ 0.72 12.35 $\pm$ 0.82 14.37 $\pm$ 0.80
en-es -0.97 $\pm$ 0.61 6.48 $\pm$ 0.69 7.45 $\pm$ 0.69
en-fr 0.04 $\pm$ 0.82 10.40 $\pm$ 0.91 10.36 $\pm$ 0.91
en-it 0.20 $\pm$ 0.55 11.09 $\pm$ 0.67 10.89 $\pm$ 0.61
en-ko -0.20 $\pm$ 0.62 14.27 $\pm$ 0.57 14.47 $\pm$ 0.53
en-nl -0.44 $\pm$ 0.67 11.50 $\pm$ 0.74 11.94 $\pm$ 0.74
en-pt -6.49 $\pm$ 0.71 7.78 $\pm$ 0.86 14.28 $\pm$ 0.80
en-ru -1.36 $\pm$ 0.78 15.33 $\pm$ 0.80 16.69 $\pm$ 0.78
es-en -3.99 $\pm$ 0.79 5.45 $\pm$ 0.95 9.43 $\pm$ 0.79
fr-en -3.74 $\pm$ 0.88 7.09 $\pm$ 0.92 10.83 $\pm$ 0.93
it-en -0.79 $\pm$ 0.76 7.03 $\pm$ 0.94 7.82 $\pm$ 0.95
ko-en -4.61 $\pm$ 0.63 5.52 $\pm$ 0.68 10.13 $\pm$ 0.70
nl-en -3.74 $\pm$ 0.74 5.37 $\pm$ 1.03 9.11 $\pm$ 1.03
pt-en -6.04 $\pm$ 0.73 9.46 $\pm$ 1.15 15.50 $\pm$ 1.07
ru-en -4.11 $\pm$ 0.69 5.46 $\pm$ 1.05 9.58 $\pm$ 0.99
COMET
de-en 0.92 $\pm$ 0.50 6.60 $\pm$ 0.79 5.68 $\pm$ 0.86
en-de 1.19 $\pm$ 0.57 16.89 $\pm$ 0.93 15.70 $\pm$ 0.89
en-es 1.19 $\pm$ 0.59 10.46 $\pm$ 0.92 9.27 $\pm$ 0.89
en-fr 0.35 $\pm$ 0.48 10.86 $\pm$ 0.71 10.50 $\pm$ 0.73
en-it 0.83 $\pm$ 0.45 14.74 $\pm$ 0.78 13.90 $\pm$ 0.77
en-ko 5.16 $\pm$ 0.73 24.62 $\pm$ 0.94 19.45 $\pm$ 0.89
en-nl 1.26 $\pm$ 0.53 19.53 $\pm$ 0.78 18.27 $\pm$ 0.83
en-pt -1.85 $\pm$ 0.47 10.32 $\pm$ 0.86 12.17 $\pm$ 0.85
en-ru -0.53 $\pm$ 0.54 17.45 $\pm$ 0.87 17.98 $\pm$ 0.85
es-en 0.68 $\pm$ 0.62 7.64 $\pm$ 0.81 6.97 $\pm$ 0.84
fr-en 0.31 $\pm$ 0.55 7.79 $\pm$ 0.83 7.48 $\pm$ 0.84
it-en 2.57 $\pm$ 0.72 7.76 $\pm$ 0.71 5.19 $\pm$ 0.89
ko-en -0.78 $\pm$ 0.41 6.51 $\pm$ 0.59 7.29 $\pm$ 0.61
nl-en -0.68 $\pm$ 0.57 7.58 $\pm$ 0.89 8.26 $\pm$ 0.95
pt-en -1.81 $\pm$ 0.40 9.40 $\pm$ 1.07 11.21 $\pm$ 1.04
ru-en -0.57 $\pm$ 0.40 6.33 $\pm$ 0.73 6.90 $\pm$ 0.70
Table 7: Pairwise differences between models on NMT experiments, with uncertainty quantified using bootstrap resampling and multiple seeds. The error bar corresponds $\pm$ one standard error. Nonsignificant differences are indicated in boldface.
Summary variance component tables
Table 8 shows a detailed summary of variance components for NMT experimets.
BLEU
Mean SD Min Max
Between-language total ($\nu$)
aya-expanse-8b 7.11 - - -
TowerInstruct-7B 8.49 - - -
Clarus-7B 4.50 - - -
Within-language total ($\eta_{l}$)
aya-expanse-8b 0.64 0.10 0.40 0.80
TowerInstruct-7B 0.77 0.22 0.28 1.15
Clarus-7B 0.70 0.23 0.22 1.12
Seed-to-seed ($\sigma_{l}$)
aya-expanse-8b 0.34 0.10 0.15 0.55
TowerInstruct-7B 0.47 0.16 0.13 0.78
Clarus-7B 0.48 0.20 0.14 0.92
Boot-to-boot ($\tau_{l}$)
aya-expanse-8b 0.53 0.08 0.34 0.67
TowerInstruct-7B 0.60 0.17 0.25 0.93
Clarus-7B 0.49 0.13 0.17 0.70
ChrF++
Between-language total ($\nu$)
aya-expanse-8b 8.05 - - -
TowerInstruct-7B 8.99 - - -
Clarus-7B 9.66 - - -
Within-language total ($\eta_{l}$)
aya-expanse-8b 0.53 0.06 0.43 0.68
TowerInstruct-7B 0.48 0.08 0.33 0.64
Clarus-7B 0.67 0.16 0.34 0.99
Seed-to-seed ($\sigma_{l}$)
aya-expanse-8b 0.28 0.09 0.11 0.46
TowerInstruct-7B 0.25 0.08 0.08 0.42
Clarus-7B 0.43 0.18 0.11 0.82
Boot-to-boot ($\tau_{l}$)
aya-expanse-8b 0.44 0.04 0.35 0.51
TowerInstruct-7B 0.40 0.04 0.32 0.48
Clarus-7B 0.50 0.06 0.32 0.58
COMET
Between-language total ($\nu$)
aya-expanse-8b 1.21 - - -
TowerInstruct-7B 2.44 - - -
Clarus-7B 5.75 - - -
Within-language total ($\eta_{l}$)
aya-expanse-8b 0.36 0.07 0.26 0.55
TowerInstruct-7B 0.38 0.10 0.23 0.63
Clarus-7B 0.74 0.12 0.53 1.01
Seed-to-seed ($\sigma_{l}$)
aya-expanse-8b 0.20 0.06 0.10 0.34
TowerInstruct-7B 0.22 0.11 0.08 0.52
Clarus-7B 0.50 0.15 0.27 0.87
Boot-to-boot ($\tau_{l}$)
aya-expanse-8b 0.30 0.06 0.22 0.46
TowerInstruct-7B 0.30 0.06 0.21 0.44
Clarus-7B 0.53 0.06 0.41 0.66
Table 8: Summary of variance components for NMT experiments.
Mean SD Min Max
Between-language total ($\sigma$)
Glot500 (concat) 6.23 - - -
Glot500 (ind) 8.16 - - -
mBERT (concat) 14.28 - - -
mBERT (ind) 9.83 - - -
XLM-R (concat) 6.24 - - -
XLM-R (ind) 7.04 - - -
Within-language total ($\eta_{l}$)
Glot500 (concat) 1.76 0.93 0.51 4.24
Glot500 (ind) 1.49 0.52 0.68 2.74
mBERT (concat) 1.27 0.76 0.12 3.14
mBERT (ind) 1.26 0.75 0.53 3.53
XLM-R (concat) 1.58 1.24 0.48 5.45
XLM-R (ind) 1.47 1.11 0.37 4.80
Seed-to-seed ($\sigma_{l}$)
Glot500 (concat) 0.84 0.62 0.13 2.63
Glot500 (ind) 0.75 0.41 0.23 1.85
mBERT (concat) 0.71 0.57 0.04 2.34
mBERT (ind) 0.66 0.48 0.19 2.10
XLM-R (concat) 0.64 0.56 0.15 2.40
XLM-R (ind) 0.66 0.51 0.14 2.18
Boot-to-boot ($\tau_{l}$)
Glot500 (concat) 1.54 0.81 0.47 3.72
Glot500 (ind) 1.28 0.40 0.59 2.18
mBERT (concat) 1.02 0.67 0.12 2.79
mBERT (ind) 1.05 0.60 0.46 2.87
XLM-R (concat) 1.13 1.09 0.40 4.77
XLM-R (ind) 1.30 0.74 0.33 3.28
Table 9: Summary of variance components for NER experiments.
Detailed variance component tables
Table 10 shows detailed estimates for variance components for QA. Table 11 shows detailed estimates for variance components for MT. Table 12 shows detailed estimates for variance components for NER.
Clarus-7B TowerInstruct-7B aya-expanse-8b gemma-2-9b
Language $\sigma$ $\tau$ $\eta$ $\sigma$ $\tau$ $\eta$ $\sigma$ $\tau$ $\eta$ $\sigma$ $\tau$ $\eta$
Arabic 0.37 0.66 0.76 0.57 0.60 0.83 0.27 1.08 1.11 0.91 0.92 1.29
German 0.35 0.64 0.73 0.83 0.86 1.20 0.51 1.13 1.24 0.60 1.10 1.25
Greek 0.34 0.84 0.91 0.42 0.51 0.66 1.00 1.02 1.43 0.82 1.01 1.30
English 0.65 1.13 1.30 0.31 0.98 1.02 0.86 1.01 1.33 0.85 1.17 1.44
Spanish 1.22 1.06 1.62 0.91 0.85 1.24 0.47 1.06 1.16 1.21 0.97 1.55
Hindi 0.14 0.83 0.84 0.20 0.37 0.42 0.59 1.13 1.28 0.63 1.03 1.21
Romanian 0.38 0.83 0.91 0.76 0.84 1.13 0.60 1.10 1.25 1.17 1.06 1.58
Russian 0.75 0.83 1.12 0.49 1.00 1.11 0.45 1.12 1.21 1.20 0.95 1.53
Thai 0.69 0.89 1.12 0.24 0.45 0.51 0.90 1.13 1.44 0.70 1.01 1.23
Turkish 0.18 0.57 0.60 0.27 0.60 0.66 0.77 1.11 1.35 0.63 1.07 1.24
Vietnamese 0.59 0.50 0.77 0.31 0.65 0.72 0.57 1.03 1.18 0.68 0.98 1.20
Chinese 0.57 0.70 0.90 0.88 0.80 1.19 0.56 1.27 1.39 0.56 1.02 1.16
Table 10: Detailed estimates of the standard deviations due to model-related $\sigma$ and bootstrap-related $\tau$ uncertainty in question answering experiments.
aya-expanse-8b TowerInstruct-7B Clarus-7B
$\sigma$ $\tau$ $\sigma$ $\tau$ $\sigma$ $\tau$
BLEU
German - English 0.39 0.57 0.67 0.93 0.58 0.70
English - German 0.45 0.54 0.59 0.65 0.24 0.48
English - Spanish 0.17 0.41 0.48 0.44 0.44 0.44
English - French 0.45 0.66 0.62 0.64 0.82 0.67
English - Italian 0.24 0.47 0.18 0.49 0.46 0.41
English - Korean 0.20 0.34 0.13 0.25 0.14 0.17
English - Dutch 0.37 0.45 0.48 0.49 0.22 0.37
English - Portuguese 0.23 0.67 0.40 0.65 0.92 0.65
English - Russian 0.38 0.52 0.41 0.55 0.15 0.41
Spanish - English 0.55 0.51 0.43 0.61 0.36 0.43
French - English 0.43 0.63 0.78 0.76 0.58 0.57
Italian - English 0.26 0.53 0.65 0.66 0.59 0.51
Korean - English 0.30 0.50 0.51 0.57 0.34 0.47
Dutch - English 0.15 0.55 0.59 0.60 0.52 0.47
Portuguese - English 0.46 0.61 0.32 0.68 0.79 0.59
Russian - English 0.38 0.55 0.25 0.60 0.58 0.57
ChrF++
German - English 0.32 0.43 0.33 0.43 0.51 0.58
English - German 0.30 0.43 0.24 0.42 0.38 0.51
English - Spanish 0.24 0.36 0.29 0.32 0.35 0.41
English - French 0.33 0.47 0.39 0.44 0.47 0.54
English - Italian 0.25 0.35 0.08 0.32 0.31 0.41
English - Korean 0.26 0.38 0.18 0.38 0.11 0.32
English - Dutch 0.28 0.37 0.30 0.37 0.39 0.42
English - Portuguese 0.21 0.50 0.21 0.41 0.39 0.53
English - Russian 0.31 0.45 0.34 0.41 0.19 0.53
Spanish - English 0.46 0.50 0.11 0.40 0.44 0.51
French - English 0.34 0.51 0.42 0.48 0.40 0.56
Italian - English 0.34 0.40 0.34 0.42 0.57 0.53
Korean - English 0.14 0.41 0.20 0.41 0.16 0.50
Dutch - English 0.11 0.51 0.29 0.44 0.72 0.52
Portuguese - English 0.33 0.49 0.11 0.40 0.82 0.56
Russian - English 0.29 0.46 0.17 0.41 0.70 0.54
COMET
German - English 0.14 0.22 0.32 0.28 0.60 0.44
English - German 0.31 0.33 0.13 0.32 0.55 0.61
English - Spanish 0.34 0.28 0.26 0.29 0.55 0.57
English - French 0.14 0.29 0.22 0.29 0.27 0.57
English - Italian 0.10 0.31 0.12 0.28 0.45 0.54
English - Korean 0.31 0.46 0.18 0.44 0.42 0.62
English - Dutch 0.15 0.29 0.24 0.34 0.42 0.57
English - Portuguese 0.19 0.30 0.21 0.23 0.59 0.54
English - Russian 0.25 0.33 0.17 0.30 0.40 0.66
Spanish - English 0.24 0.33 0.25 0.38 0.49 0.50
French - English 0.19 0.31 0.22 0.34 0.55 0.51
Italian - English 0.23 0.25 0.52 0.36 0.39 0.49
Korean - English 0.15 0.22 0.18 0.25 0.34 0.41
Dutch - English 0.11 0.30 0.36 0.30 0.69 0.46
Portuguese - English 0.20 0.26 0.08 0.21 0.87 0.51
Russian - English 0.19 0.25 0.11 0.22 0.48 0.45
Table 11: Detailed estimates of the standard deviations due to model-related $\sigma$ and bootstrap-related $\tau$ uncertainty in machine translation experiments.
Glot500 mBERT XLM-R
Concat Individual Concat Individual Concat Individual
$\sigma$ $\tau$ $\sigma$ $\tau$ $\sigma$ $\tau$ $\sigma$ $\tau$ $\sigma$ $\tau$ $\sigma$ $\tau$
$\text{ara}_{\text{AQMAR}}$ 0.61 1.61 0.61 1.87 1.60 2.70 2.05 2.87 0.32 0.72 0.45 1.05
$\text{bam}_{\text{MasakhaNER\_2.0}}$ 0.77 1.48 0.79 1.98 1.46 2.11 1.06 2.02 0.27 0.73 0.19 0.33
$\text{bar}_{\text{BarNER}}$ 1.28 2.08 1.85 2.02 1.57 2.09 2.10 2.25 0.52 0.94 0.18 0.34
$\text{bbj}_{\text{MasakhaNER\_2.0}}$ 0.75 1.73 1.59 1.79 0.89 1.94 0.49 0.83 0.68 0.93 0.23 0.45
$\text{cat}_{\text{AnCora}}$ 0.82 1.72 0.81 1.67 0.29 0.12 0.58 0.84 0.52 1.22 0.37 0.45
$\text{cmn}_{\text{UNER\_Chinese\_GSD}}$ 1.16 1.57 0.88 1.19 1.97 0.12 0.34 0.94 0.79 1.08 0.14 0.34
$\text{cmn}_{\text{UNER\_Chinese\_GSDSIMP}}$ 0.76 1.15 0.30 1.14 0.11 0.21 0.57 0.97 0.34 0.93 0.21 0.37
$\text{dan}_{\text{DaNE}}$ 1.42 1.18 0.88 1.23 0.06 0.13 0.41 0.84 0.58 0.96 1.41 2.43
$\text{deu}_{\text{GermEval}}$ 0.95 1.35 0.59 1.27 0.07 0.13 0.59 0.92 1.98 3.06 0.61 3.07
$\text{ell}_{\text{elNER}}$ 0.84 1.26 0.74 1.17 0.04 0.12 0.41 1.06 0.61 4.24 2.18 3.14
$\text{eng}_{\text{Tweebank}}$ 0.93 1.21 0.66 1.13 0.32 0.56 0.60 1.09 1.20 2.96 0.61 3.28
$\text{eng}_{\text{UNER\_English\_EWT}}$ 0.83 1.32 0.78 1.49 0.27 0.53 0.67 1.34 1.11 2.50 1.07 2.41
$\text{eus}_{\text{EIEC}}$ 1.54 2.81 0.92 1.44 0.41 0.74 0.56 1.16 2.40 2.66 1.86 2.45
$\text{ewe}_{\text{MasakhaNER\_2.0}}$ 2.63 3.29 0.71 1.93 0.38 0.69 0.27 1.12 0.93 3.08 0.26 0.56
$\text{fin}_{\text{TurkuNLP}}$ 2.10 3.68 0.83 1.69 0.24 0.57 0.47 1.08 0.29 0.52 0.22 0.51
$\text{fon}_{\text{MasakhaNER\_2.0}}$ 1.47 3.72 0.85 1.41 0.39 0.56 0.89 1.41 0.35 0.49 0.28 0.63
$\text{glg}_{\text{SLI\_Galician\_Corpora}}$ 1.29 2.80 0.89 1.47 0.65 0.88 1.30 1.43 0.40 0.66 0.25 0.69
$\text{hau}_{\text{MasakhaNER}}$ 2.02 3.27 0.42 0.59 1.40 0.95 0.77 1.67 0.29 0.54 0.24 0.55
$\text{hau}_{\text{MasakhaNER\_2.0}}$ 0.80 1.56 0.46 0.67 0.72 1.03 1.24 1.52 0.18 0.48 0.23 0.52
$\text{heb}_{\text{NEMO\_SPMRL}}$ 1.04 1.64 0.54 0.70 0.79 1.29 0.55 1.36 0.19 0.47 0.56 0.80
$\text{heb}_{\text{NEMO\_UD}}$ 1.23 1.55 0.43 0.79 0.60 0.93 1.21 1.46 0.72 1.29 0.77 0.86
$\text{hin}_{\text{HiNER}}$ 1.15 1.54 0.32 0.60 0.96 0.97 0.27 0.54 1.64 1.64 0.64 0.96
$\text{hrv}_{\text{hr500k}}$ 0.92 1.62 0.45 0.70 0.20 0.43 0.27 0.56 0.96 1.74 1.26 1.09
$\text{ibo}_{\text{MasakhaNER}}$ 0.90 1.76 0.76 1.27 0.91 0.52 0.40 1.03 1.67 1.68 0.66 0.87
$\text{ibo}_{\text{MasakhaNER\_2.0}}$ 0.24 0.72 1.13 1.34 0.60 0.63 0.36 0.72 0.72 1.26 0.84 0.93
$\text{kaz}_{\text{KazNERD}}$ 0.67 0.72 0.68 1.60 0.55 0.79 0.31 0.55 1.19 1.56 0.62 1.35
$\text{kin}_{\text{MasakhaNER}}$ 0.41 0.99 0.83 1.43 0.22 0.46 0.36 0.58 0.35 0.73 0.55 1.51
$\text{kin}_{\text{MasakhaNER\_2.0}}$ 0.40 0.81 0.56 1.37 0.82 0.55 0.29 0.58 0.42 0.80 0.91 1.74
$\text{lug}_{\text{MasakhaNER}}$ 0.46 0.72 0.85 1.23 0.26 0.61 0.41 0.61 0.54 1.16 0.75 1.80
$\text{lug}_{\text{MasakhaNER\_2.0}}$ 0.34 0.67 0.68 1.17 2.34 0.46 0.93 1.03 0.59 0.89 0.44 1.47
$\text{luo}_{\text{MasakhaNER}}$ 0.86 1.40 0.70 1.30 0.57 1.06 0.40 0.71 0.53 0.69 1.08 1.58
$\text{luo}_{\text{MasakhaNER\_2.0}}$ 0.61 1.56 0.92 1.38 0.39 0.50 0.19 0.61 0.44 0.65 0.83 2.52
$\text{mar}_{\text{L3Cube\_MahaNER}}$ 1.08 2.70 1.05 1.37 0.37 0.60 0.34 0.58 0.84 4.77 0.61 2.70
$\text{mos}_{\text{MasakhaNER\_2.0}}$ 0.46 1.42 0.74 1.29 0.61 0.42 0.39 1.14 1.14 1.07 1.19 2.67
$\text{nep}_{\text{EverestNER}}$ 0.78 1.44 0.47 1.29 0.73 1.31 0.54 1.13 0.91 1.07 1.22 2.85
$\text{nld}_{\text{CONLL02}}$ 0.56 1.61 0.75 1.02 0.88 1.57 1.28 1.65 1.71 1.07 1.11 2.48
$\text{nno}_{\text{NorNE}}$ 0.79 1.56 0.56 1.08 1.04 1.56 0.68 1.42 0.93 1.07 1.42 2.76
$\text{nob}_{\text{NorNE}}$ 0.59 1.61 0.81 1.18 2.06 1.68 0.46 1.15 1.57 1.07 0.28 0.83
$\text{nya}_{\text{MasakhaNER\_2.0}}$ 1.02 3.07 0.61 1.14 0.57 1.37 0.79 1.13 0.26 0.40 0.56 0.94
$\text{pcm}_{\text{MasakhaNER}}$ 0.47 1.52 0.75 0.99 1.21 1.47 0.32 0.95 0.26 0.44 0.51 1.03
$\text{pcm}_{\text{MasakhaNER\_2.0}}$ 0.65 1.56 0.38 1.06 0.30 0.57 0.80 1.02 0.39 0.55 0.57 1.05
$\text{por}_{\text{UNER Portuguese}}$ 0.59 1.62 0.23 1.08 0.48 0.67 1.57 1.52 0.49 0.68 0.44 0.99
$\text{qaf}_{\text{UNER Arabizi}}$ 1.22 2.00 0.43 1.02 0.32 0.71 0.97 1.31 0.33 0.40 1.10 1.06
$\text{ron}_{\text{RONEC}}$ 1.47 1.99 0.87 1.26 0.49 0.90 0.53 0.92 0.48 0.64 0.32 0.61
$\text{slk}_{\text{UNER\_Slovak\_SNK}}$ 1.50 2.36 0.49 1.22 0.25 0.64 0.68 0.99 0.33 0.54 0.32 0.65
$\text{slk}_{\text{WikiGoldSK}}$ 1.05 1.98 0.52 1.08 0.71 0.71 0.38 0.47 0.15 0.47 0.43 0.84
$\text{slv}_{\text{ssj500k}}$ 1.17 2.02 0.61 0.91 0.65 1.85 0.27 0.46 0.34 0.64 0.43 0.94
$\text{sna}_{\text{MasakhaNER\_2.0}}$ 0.65 1.71 0.48 0.64 1.68 2.02 0.47 0.60 0.26 0.55 0.38 0.62
$\text{spa}_{\text{AnCora}}$ 0.18 0.55 0.69 0.67 0.68 1.98 0.65 0.65 0.22 0.53 0.63 0.68
$\text{spa}_{\text{CONLL02}}$ 0.18 0.47 0.65 0.86 1.21 1.81 0.19 0.51 0.22 0.46 0.44 1.62
$\text{swa}_{\text{MasakhaNER}}$ 0.38 0.67 1.07 0.97 0.65 1.83 0.33 0.53 0.24 0.68 0.61 1.62
$\text{swa}_{\text{MasakhaNER\_2.0}}$ 0.30 0.52 0.46 0.65 1.29 1.97 0.37 0.97 0.43 0.68 0.50 1.66
$\text{swe}_{\text{UNER\_Swedish\_Talkbanken}}$ 0.22 0.53 0.61 0.66 0.31 0.56 0.86 1.15 0.28 0.70 0.58 1.48
$\text{tha}_{\text{ThaiNNER}}$ 0.29 0.50 0.57 1.42 0.22 0.57 0.67 1.25 0.19 0.70 0.64 1.71
$\text{tsn}_{\text{MasakhaNER\_2.0}}$ 0.29 0.67 1.29 1.86 0.48 0.66 0.96 1.51 0.50 0.68 1.11 1.61
$\text{twi}_{\text{MasakhaNER\_2.0}}$ 0.24 0.62 0.73 2.18 0.65 0.74 0.62 1.11 0.38 0.63 0.43 0.79
$\text{wol}_{\text{MasakhaNER}}$ 0.45 0.88 1.63 1.75 0.41 0.56 1.70 1.33 0.45 0.91 0.86 0.79
$\text{wol}_{\text{MasakhaNER\_2.0}}$ 0.37 0.72 0.73 1.43 0.29 0.64 0.37 0.67 0.48 1.09 0.75 1.02
$\text{xho}_{\text{MasakhaNER\_2.0}}$ 0.27 0.64 0.68 1.68 0.77 2.16 0.50 0.69 0.44 1.12 0.57 0.88
$\text{yor}_{\text{MasakhaNER}}$ 0.13 0.66 1.12 2.14 0.61 2.79 0.65 0.88 0.68 1.22 0.68 0.81
$\text{yor}_{\text{MasakhaNER\_2.0}}$ 1.63 1.75 1.60 2.15 1.25 2.24 0.75 0.81 0.49 0.91 0.52 0.82
Table 12: Detailed estimates of the standard deviations due to model-related $\sigma$ and bootstrap-related $\tau$ uncertainty in NER experiments.
C.2 Aggregate inferences
Table 13 shows uncertainty quantified estimates of aggregate performance for the NER experiments.
Model Statistic F1 (true) Estimate SE 95% CI (percentile) 95% CI ($\hat{\theta}\pm$ 2SE) CI ($\hat{\theta}\pm$ IQW)
Glot500 (c.) $\hat{\mu}_{\text{am}}$ 85.62 85.61 0.20 85.21 85.99 85.21 86.01 85.22 86.00
$\hat{\mu}_{\text{gm}}$ 85.62 85.35 0.22 84.90 85.77 84.91 85.79 84.92 85.79
$\mu_{\text{md}}$ 85.62 87.53 0.38 86.73 88.23 86.77 88.29 86.78 88.28
Glot500 (i.) $\hat{\mu}_{\text{am}}$ 83.71 83.71 0.41 82.82 84.40 82.89 84.52 82.91 84.50
$\hat{\mu}_{\text{gm}}$ 83.71 83.17 0.55 81.90 84.04 82.07 84.27 82.10 84.24
$\mu_{\text{md}}$ 83.71 86.46 0.42 85.63 87.28 85.62 87.30 85.64 87.29
mBERT (c.) $\hat{\mu}_{\text{am}}$ 72.60 72.61 0.24 72.13 73.07 72.12 73.09 72.13 73.08
$\hat{\mu}_{\text{gm}}$ 72.60 70.57 0.30 69.97 71.16 69.96 71.18 69.98 71.17
$\mu_{\text{md}}$ 72.60 76.34 0.49 75.36 77.24 75.37 77.31 75.40 77.28
mBERT (i.) $\hat{\mu}_{\text{am}}$ 77.31 77.31 0.40 76.40 77.92 76.52 78.11 76.56 78.07
$\hat{\mu}_{\text{gm}}$ 77.31 76.51 0.58 75.10 77.27 75.35 77.67 75.42 77.60
$\mu_{\text{md}}$ 77.31 79.61 0.48 78.70 80.53 78.65 80.57 78.69 80.52
XLM-R (c.) $\hat{\mu}_{\text{am}}$ 84.91 84.88 0.20 84.47 85.28 84.47 85.29 84.48 85.29
$\hat{\mu}_{\text{gm}}$ 84.91 84.62 0.23 84.17 85.06 84.16 85.07 84.17 85.06
$\mu_{\text{md}}$ 84.91 86.71 0.39 85.95 87.47 85.94 87.49 85.95 87.47
XLM-R (i.) $\hat{\mu}_{\text{am}}$ 83.88 83.87 0.49 82.41 84.43 82.88 84.85 82.85 84.88
$\hat{\mu}_{\text{gm}}$ 83.88 83.28 1.19 79.59 84.14 80.89 85.66 81.01 85.55
$\mu_{\text{md}}$ 83.88 86.17 0.52 85.09 87.14 85.13 87.22 85.15 87.20
Glot500 (c.) $\hat{\mu}_{\text{am}}$ 85.62 85.85 0.58 84.72 87.01 84.69 87.01 84.70 86.99
$\hat{\mu}_{\text{gm}}$ 85.62 85.60 0.62 84.39 86.85 84.37 86.84 84.37 86.83
$\mu_{\text{md}}$ 85.62 87.67 0.66 86.16 88.75 86.35 88.99 86.37 88.96
Glot500 (i.) $\hat{\mu}_{\text{am}}$ 83.71 83.81 0.84 82.12 85.39 82.14 85.49 82.18 85.44
$\hat{\mu}_{\text{gm}}$ 83.71 83.27 1.01 81.10 85.07 81.25 85.29 81.28 85.25
$\mu_{\text{md}}$ 83.71 86.55 0.74 84.96 87.89 85.06 88.04 85.08 88.01
mBERT (c.) $\hat{\mu}_{\text{am}}$ 72.60 73.58 1.30 71.07 76.14 70.97 76.18 71.04 76.12
$\hat{\mu}_{\text{gm}}$ 72.60 71.67 1.71 68.47 75.18 68.25 75.10 68.32 75.03
$\mu_{\text{md}}$ 72.60 76.99 0.95 75.12 78.95 75.10 78.89 75.08 78.91
mBERT (i.) $\hat{\mu}_{\text{am}}$ 77.31 77.24 0.93 75.32 79.05 75.38 79.10 75.37 79.11
$\hat{\mu}_{\text{gm}}$ 77.31 76.45 1.10 74.07 78.52 74.25 78.65 74.23 78.67
$\mu_{\text{md}}$ 77.31 79.57 0.80 78.03 81.16 77.96 81.17 78.00 81.13
XLM-R (c.) $\hat{\mu}_{\text{am}}$ 84.91 85.05 0.57 83.95 86.16 83.91 86.18 83.94 86.15
$\hat{\mu}_{\text{gm}}$ 84.91 84.79 0.60 83.63 85.99 83.58 86.00 83.61 85.97
$\mu_{\text{md}}$ 84.91 86.79 0.68 85.40 88.06 85.43 88.16 85.46 88.12
XLM-R (i.) $\hat{\mu}_{\text{am}}$ 83.88 83.84 0.82 81.78 85.24 82.20 85.49 82.11 85.57
$\hat{\mu}_{\text{gm}}$ 83.88 83.27 1.52 78.07 84.98 80.23 86.32 79.82 86.73
$\mu_{\text{md}}$ 83.88 86.03 0.92 84.12 87.60 84.19 87.87 84.29 87.77
Table 13: Overall model performance with languages fixed (top) and randomized (bottom).
C.3 Ranks
Table 6 shows the marginal distributions of ranks for the NER task.